OpenStudy (anonymous):

let g(x) be a horizontal compression of f(x)=4x+2 by a factor of 1/3. write the rule for g(x).

OpenStudy (anonymous):

g(x)=f(3x)=4*3x+2=12x+2

OpenStudy (anonymous):

whats the answer? simplify

OpenStudy (smurfy14):

answer is 12x+2